# MultiCrafter: High-Fidelity Multi-Subject Generation via Disentangled Attention and Identity-Aware Preference Alignment

### v1

@@ Abstract
Multi-subject image generation aims to synthesize user-provided subjects in a single image while preserving subject fidelity, ensuring prompt consistency, and aligning with human aesthetic preferences. However, existing methods, particularly those built on the In-Context-Learning paradigm, are limited by their reliance on simple reconstruction-based objectives, leading to both severe attribute leakage that compromises subject fidelity and failing to align with nuanced human preferences. To address this, we propose MultiCrafter, a framework that ensures high-fidelity, preference-aligned generation. First, we find that the root cause of attribute leakage is a significant entanglement of attention between different subjects during the generation process. Therefore, we introduce explicit positional supervision to explicitly separate attention regions for each subject, effectively mitigating attribute leakage. To enable the model to accurately plan the attention region of different subjects in diverse scenarios, we employ a Mixture-of-Experts architecture to enhance the model’s capacity, allowing different experts to focus on different scenarios. Finally, we design a novel online reinforcement learning framework to align the model with human preferences, featuring a scoring mechanism to accurately assess multi-subject fidelity and a more stable training strategy tailored for the MoE architecture. Experiments validate that our framework significantly improves subject fidelity while aligning with human preferences better.
Figure 1: MultiCrafter enables multi-subject personalization. Inputs are surrounded by squares.
@@ 1 Introduction
Subject-driven image generation, which aims to create images featuring user-provided subjects, has become a cornerstone of personalized content creation. Propelled by higher-quality data and the widespread adoption of Diffusion Transformers (Labs, 2024; Esser et al., 2024; Gao et al., 2025; Li et al., 2024b), text-to-image models have seen rapid advancements. This progress has significantly enhanced single-subject image generation, where models now excel at preserving subject fidelity (Li et al., 2025d; Feng et al., 2025; He et al., 2025). Among the various techniques, the In-Context-Learning (ICL) paradigm (Huang et al., 2024a; b; c; Tan et al., 2024; Liang et al., 2025) has emerged as a mainstream approach. Its high adaptability to the transformer architecture and strong capability for maintaining subject fidelity have made it particularly effective for single-subject tasks.
@@
However, extending this success from single-subject to multi-subject generation is not trivial. Building on the ICL framework, recent works like UNO (Wu et al., 2025c) and OmniGen (Xiao et al., 2025) have ventured into this complex multi-subject customization task. Despite their efforts, these methods often produce suboptimal results, especially when dealing with intricate subjects like human faces. They frequently suffer from suboptimal results, such as identity fusion and attribute leakage between subjects. This raises a critical question that forms the very foundation of our work: why do existing ICL-based methods falter in the multi-subject setting?
The primary challenge stems from the training objective. Existing ICL-based methods are typically optimized with a simple reconstruction loss. This objective implicitly tasks the model with the dual responsibilities of distinguishing subject features and arranging them spatially. However, such supervision alone proves insufficient for the complexities of multi-subject scenarios. As shown in Fig.˜2, this inadequacy leads to an undesired entanglement between subject-specific attention fields in these methods, like UNO. This phenomenon, which we term attention bleeding, causes attribute leakage and severely damages subject fidelity. Furthermore, a simple reconstruction objective fails to capture nuanced human preferences, such as aesthetic quality and precise prompt alignment.
@@
To address these limitations, we introduce MultiCrafter, a framework that achieves high-fidelity, preference-aligned multi-subject image generation through three key innovations. To address the attribute leakage caused by attention bleeding, we propose an Identity-Disentangled Attention Regularization. This mechanism applies explicit positional supervision only during the training phase to double blocks in FLUX (Labs, 2024), which are pivotal regions for feature injection and spatial control. This compels the model to distinguish between different subject features and learn distinct, disentanglement attention regions for each subject, drastically reducing attribute leakage.
Considering a single model struggles to cope with various attention layouts caused by a diverse range of subjects and prompts, we enhance its capacity by incorporating a Mixture-of-Experts (MoE) architecture. Inspired by MoE-LoRA’s success in multitask tuning (Feng et al., 2024; Zhang et al., 2025), our Efficient Adaptive Expert Tuning allows different expert networks to specialize in varied scenarios, dynamically selected via a routing mechanism. This ensure our method to maintain excellent subject fidelity in various scenarios without an increase in inference complexity.
@@
For aligning with human aesthetic and semantic preferences while ensuring subject fidelity, we design a novel online reinforcement learning framework. We introduce Identity-Preserving Preference Optimization that aligns the model across three axes: aesthetic quality, text-image alignment, and subject fidelity. To accurately measure subject fidelity, we introduce Multi-ID Alignment Reward, which use the Hungarian matching algorithm to maximize the overall match quality between multiple generated subjects and their references for precise scoring. Besides, we introduce Group Sequence Policy Optimization (GSPO) Zheng et al. (2025) to adapt to our MoE-LoRA design, thereby avoiding the core issue of training instability caused by expert-activation volatility inherent in MoE architectures. Our experiments demonstrate that MultiCrafter achieves significant improvements over existing methods. Our main contributions are as follows:
@@
  • We integrate an MoE-LoRA architecture that increases capacity for diverse subjects and spatial layouts while maintaining inference efficiency.
  • We design the first online reinforcement learning framework tailored for multi-subject generation, introducing a Multi-ID Alignment Reward and GSPO for stable and preference-aligned training.
Figure 2: Visual comparison of attention maps. The ICL-based method UNO, which only relies on reconstruction loss (left), fails to preserve the subject fidelity, where the double block’s attention regions for each subject are entangled, leading to attribute leakage. Our method overcomes this problem and maintains subject fidelity.
@@ 2 Related Work
Subject-driven Generation has attracted increasing attention with the development of diffusion models and multimodal (Chen et al., 2023; Han et al., 2023; Shi et al., 2024; Ruiz et al., 2024; Wu et al., 2024a; b; 2025d; Hua et al., 2023; Han et al., 2024; Liu et al., 2023a; b; Dou et al., 2024; Gu et al., 2024; Feng et al., 2025; Ma et al., 2025b; 2024; a; Zhao et al., 2024; Li et al., 2025c). These works can be broadly categorized into two types. (1) Fine-tuning-based methods, includes methods such as Textual Inversion (Gal et al., 2022), DreamBooth (Ruiz et al., 2023), and Custom Diffusion (Kumari et al., 2023). These approaches achieve customization by fine-tuning part of the model’s parameters. (2) Tuning-free methods, like IP-Adapter (Ye et al., 2023), InstantID (Wang et al., 2024), and PhotoMaker (Li et al., 2024a). These approaches leverage large-scale training to eliminate the need for retraining when the subject changes. However, customized generation of multiple subjects from a single image introduces new challenges, especially in maintaining individual subject fidelity and mitigating attribute entanglement. Recently, the powerful capabilities of foundation models based on the DiT (Peebles & Xie, 2023) architecture have greatly enhanced the generation of multiple subjects, leading to the emergence of a series of works such as HunyuanCustom (Hu et al., 2025), OmniControl (Tan et al., 2024), UniReal (Chen et al., 2025b), UNO (Wu et al., 2025c), DreamO (Mou et al., 2025) and XVerse (Chen et al., 2025a).
Reinforcement Learning for Text-to-Image Generation has become an active area of research. Taking advantage of the development of reinforcement learning technology Wang et al. (2025a); Pang et al. (2023); Wang et al. (2025b), initial strategies included policy gradient methods such as Proximal Policy Optimization (PPO) (Schulman et al., 2017; Black et al., 2023; Fan et al., 2023; Gupta et al., 2025; Miao et al., 2024; Zhao et al., 2025). A subsequent major development is the adoption of Direct Preference Optimization (DPO) and its variants (Wallace et al., 2024; Yang et al., 2024; Yuan et al., 2024; Liu et al., 2025b; Zhang et al., 2024; Furuta et al., 2024; Li et al., 2025a; Wang et al., 2025b). Some recent works have introduced online RL technology, Group Relative Policy Optimization (GRPO)  Shao et al. (2024), into Text-to-Image Generation, achieving significant performance gains. Flow-GRPO (Liu et al., 2025a), DanceGRPO (Xue et al., 2025) introduce exploration by reformulating the deterministic Ordinary Differential Equation (ODE) of flow-matching models into a Stochastic Differential Equation (SDE). The improved MixGRPO (Li et al., 2025b) further boosts training efficiency with a mixed ODE-SDE framework. However, these methods have been limited to the basic text-to-image task, leaving the application of online reinforcement learning to multi-subject driven generation largely unexplored.
@@ 3 Preliminary
$\mathcal{L}_{diff}=\mathbb{E}_{t,z_{0},\epsilon\in\mathcal{N}(0,1)}\|v-v_{\theta}(z_{t},t,c_{text})\|^{2},\quad v=z_{0}-\epsilon.$ (1)
The DiT framework and Flow Matching are widely used in recent diffusion models, such as Stable Diffusion3 (Esser et al., 2024) and Flux (Labs, 2024). In this paper, we use Flux as our base model.
Group Relative Policy Optimization (GRPO). GRPO struggles with ODE-based flow models because their deterministic systems lack the inherent stochasticity essential for reinforcement learning frameworks. So DanceGRPO Xue et al. (2025) and Flow-GRPO Liu et al. (2025a) convert ODEs to SDEs, enabling stochastic reinforcement learning for image generation. MixGRPO Li et al. (2025b) further improves efficiency by applying this only within a sliding time window during training. Given the prompt $c_{text}$, the training process in MixGRPO is similar to Flow-GRPO and DanceGRPO, but only optimizes the time steps sampled within the interval $S$. The behavior of this window is governed by key hyperparameters: the window size $w$ sets the number of consecutive timesteps to optimize at once, the shift interval $\tau$ determines how many training iterations pass before the window moves, and the window stride $s$ specifies how many timesteps the window advances during a shift. The final training objective is given by:
$J(\theta)=\mathbb{E}_{c\sim c_{text},\{x_{i}^{T}\}_{i=0}^{N}\sim\pi_{\theta_{\text{old}}}(\cdot|c)}\Bigg[\frac{1}{N}\sum_{i=1}^{N}\frac{1}{|S|}\sum_{t\in S}\min\Big(r_{i}^{t}(\theta)A_{i},clip(r_{i}^{t}(\theta),1-\beta,1+\beta)A_{i}\Big)\Bigg],$ (2)
where $r_{i}^{t}(\theta)$ is the policy ratio and $A_{i}$ is the advantage score. $\beta$ is a hyperparameter that serves to clip the policy ratio, ensuring stable updates, and
$r_{i}^{t}(\theta)=\frac{\pi_{\theta}(x_{t+1}\mid x_{t},c)}{\pi_{\theta_{\text{old}}}(x_{t+1}\mid x_{t},c)},\quad A_{i}=\frac{R(x_{i}^{T},c)-mean\{R(x_{i}^{T},c)\}_{i=1}^{N}}{std\{R(x_{i}^{T},c)\}_{i=1}^{N}},$ (3)
where $\pi_{\theta}(x_{t+1}\mid x_{t},c)$ is the policy function and $R(x_{i}^{0},c)$ is provided by the reward model.
@@ 4 Methods
Figure 3: Overall pipeline of MultiCrafter. Our framework is built on three core innovations: (Top Left) Identity-Disentangled Attention Regularization uses positional supervision to prevent attribute leakage; (Top Right) the MoE-LORA architecture boosts model capacity for diverse scenarios; and (Bottom) the Identity-Preserving Preference Alignment framework employs a novel online reinforcement learning strategy with a Multi-ID Alignment Reward and the stable GSPO algorithm to align the model with human preferences.
In this section, we first explore the underlying causes of feature entanglement in multi-subject customized generation and present the overall framework of our method in Sec.˜4.1. Then, in Sec.˜4.2, we introduce Identity-Disentangled Attention Regularization to address attribute leakage issues arising from attention bleeding. To enhance the model’s ability to maintain subject fidelity in different scenarios, we introduce Efficient Adaptive Expert Tuning in Sec.˜4.3. Lastly, we incorporate Identity-Preserving Preference Optimization in Sec.˜4.4, leveraging reinforcement learning to align the model with human preferences.
4.1 Explore the In-Cotext-Learning
[…]
As shown on the left of Fig.˜2, we identify two key phenomena. First, the double blocks of Flux are far more pivotal in determining the spatial layout of the reference subject than the single block. Second, methods trained solely on a reconstruction loss lead to an undesired entanglement between subject-specific attention fields, causing attribute leakage and severely compromising subject fidelity. We can therefore infer a critical condition for high fidelity: the peak response within the double block’s attention scores, corresponding to a specific subject, must consistently align with that subject’s spatial region in the generated output. We leverage this in our method. Furthermore, this simple supervision method fails to account for human preferences adequately. Therefore, we introduce reinforcement learning as a post-training to align human preferences.
@@ 4.2 Identity-Disentangled Attention Regularization
Based on that, the subject fidelity can be enhanced by strictly aligning the hot spot areas of the attention scores of double blocks with the spatial positions of the corresponding subjects in the generated image. We design a simple but effective regularization. Specifically, during the training process, for the latent space feature $\mathbf{z}_{i}^{ref}$ of the $i$-thth reference subject. Following (Wu et al., 2025c), we partition the reference feature $\mathbf{z}_{i}^{ref}$ into patches and apply positional encoding, resulting in a sequence of 1D tokens $\mathbf{z}_{i}^{r\prime}\in\mathbb{R}^{l\times c}$, $l$ is the number of tokens, $c$ is the number of channels. Then, we can obtain its attention map with the generated content $z_{t}$ at the $k$-th double block:
@@
where $\mathbf{Q}_{k,i}\in\mathbb{R}^{l\times c}$ is the query generated from the $i$-th subject image within the $k$-th double block, $\mathbf{K}\in\mathbb{R}^{l_{t}\times c}$ is the key produced by the noisy image latent tokens in the current layer, and $l_{t}$ denotes the number of tokens in the noisy image latent. For a model with $K$ double blocks, we can obtain the attention maps corresponding to the $i$-th subject from all blocks, which are aggregated into a set $\{m_{1}^{i},m_{2}^{i},...,m_{K}^{i}\}$. We then average and normalize this set to obtain the mean attention map $\hat{M}_{i}$. By pre-annotating the training data, we obtain the ground-truth mask ${M}_{i}$ corresponding to the $i$-th subject within the generated image. Notably, for human subjects, we exclusively use the facial region as the reference image and similarly focus only on the facial area for the generated image.
Finally, we employ the dice loss (Milletari et al., 2016), a standard loss function for segmentation tasks, to minimize the discrepancy between each ground-truth mask ${M}_{i}$ and the corresponding mean attention map $\hat{M}_{i}$. The formulation is as follows:
@@
4.3 Efficient Adaptive Expert Tuning
For efficient training, existing ICL-based methods usually use LoRA to fine-tune the model. But the significant variance in the spatial layout of subjects across various subjects and prompts poses a challenge for a standard LoRA, whose limited capacity is insufficient for diverse subject-driven generation. To address this, and inspired by the success of MoE-LoRA in multitask tuning (Feng et al., 2024; Liu et al., 2024; Gou et al., 2023), we adopt an MoE-LoRA architecture to expand model capacity without a substantial increase in inference overhead. This enables different experts to focus on spatial layout for a variety of scenarios, effectively tackling the challenge of scene diversity.
We strategically integrate the MoE-LoRA into the output feed-forward network (FFN) layers of the Flux, while other layers are adapted using standard LoRA for parameter efficiency. Specifically, given an input vector $h$ to the FFN layer, we define the number of experts as $N_{e}$, the rank of each LoRA as $r$, and the scaling factor as $\alpha$. A lightweight gating network, $g_{\theta}$, which dynamically routes the input vector $h$ to the most suitable experts, computes a vector of logits, $p\in\mathbb{R}^{N_{e}}$, for the experts:
$p=\operatorname{Softmax}(\operatorname{TopK}(W_{g}\cdot h,k))$
@@ 4.3 Efficient Adaptive Expert Tuning
$h_{out}=\text{FFN}(h)+\sum_{i=1}^{N_{e}}p_{i}\cdot\left(\frac{\alpha}{r}\cdot W_{B}^{i}\cdot W_{A}^{i}\cdot h\right).$

### v2

@@ Abstract
Multi-subject image generation aims to synthesize user-provided subjects in a single image while preserving subject fidelity, ensuring prompt consistency, and aligning with human aesthetic preferences. Existing In-Context-Learning based methods are limited by their highly coupled training paradigm. These methods attempt to achieve both high subject fidelity and multi-dimensional human preference alignment within a single training stage, relying on a single, indirect reconstruction loss, which is difficult to simultaneously satisfy both these goals. To address this, we propose MultiCrafter, a framework that decouples this task into two distinct training stages. First, in a pre-training stage, we introduce an explicit positional supervision mechanism that effectively resolves attention bleeding and drastically enhances subject fidelity. Second, in a post-training stage, we propose Identity-Preserving Preference Optimization, a novel online reinforcement learning framework. We feature a scoring mechanism to accurately assess multi-subject fidelity based on the Hungarian matching algorithm, which allows the model to optimize for aesthetics and prompt alignment while ensuring subject fidelity achieved in the first stage. Experiments validate that our decoupling framework significantly improves subject fidelity while aligning with human preferences better.
Figure 1: MultiCrafter enables high-quality multi-subject personalization. Inputs are surrounded by squares.
@@ 1 Introduction
With improvements in data quality and the large-scale adoption of Diffusion Transformers (DiT) [27, 8, gong2025seedream2, 14, 32], text-to-image models have developed rapidly, driving demand for personalized generation. Multi-subject image generation, which aims to create images featuring multiple user-provided subjects, is a challenging but significant task for personalized generation. This task must simultaneously satisfy two distinct and challenging objectives: (i) achieving high subject fidelity for multiple subjects, and (ii) aligning with human preferences, encompassing aesthetic quality, semantic accuracy, and precise text alignment.
@@
Mainstream In-Context-Learning (ICL) [24, 25, 53] based works, such as UNO [59] and OmniGen [62, 58], have demonstrated a high degree of adaptability to the transformer architecture, significantly improving the quality of multi-subject generation. However, despite their efforts, these methods frequently suffer from suboptimal results, such as low subject fidelity caused by attribute leakage between subjects. This raises a critical question: why do existing ICL-based methods falter in the multi-subject setting?
It is difficult to achieve these two challenging objectives simultaneously using only a highly coupled training paradigm. Specifically, these methods attempt to simultaneously achieve high subject fidelity and align with multi-dimensional human preferences, relying solely on a single and indirect reconstruction loss during a training stage. We argue that these two distinct learning objectives should not be forced into a single supervision signal, as this would inevitably force the model to find a sub-optimal result between these two objectives. This coupled learning objective produces two major failures. First, for subject fidelity, it fails to resolve the underlying spatial entanglement between subject features and their intended locations. As shown in Fig.˜2, this leads to the "attention bleeding" phenomenon (e.g., in UNO, the double block’s attention regions for each subject are entangled), causing attribute leakage and significantly harming fidelity. Second, for human preference, these methods suffer from a severe proxy-objective mismatch, as the reconstruction loss cannot directly align multi-dimensional human preferences.
@@
In this work, we introduce MultiCrafter, a multi-subject generation framework that employs a divide-and-conquer strategy to decouple this complex process by learning subject fidelity and human preference alignment separately at different training stages. We implement this via two stages focusing on different objectives: a fidelity-focused pre-training stage and a preference-optimizing post-training stage. Specifically, in the pre-training stage, we focus on achieving high subject fidelity. To address the attribute leakage caused by attention bleeding, we propose an Identity-Disentangled Attention Regularization. This mechanism applies explicit positional supervision only during the training phase to double blocks in FLUX [27], which are pivotal regions for feature injection and spatial control. During training, we add this explicit supervision to the standard reconstruction loss, guiding the model’s attention to concentrate on the precise spatial layout of each subject in the generated image. This compels the model to distinguish between different subject features and learn distinct, disentangled attention regions for each subject, drastically reducing attribute leakage. However, this regularization requires the model to master the complex and diverse spatial layouts caused by different prompts and subjects. Finding that this challenge exceeds the capacity of a single LoRA module, we use a Mixture-of-Experts LoRA (MoE-LoRA) to provide the required specialized capacity
@@
In the post-training stage, to align with multi-dimensional human preferences, we design a novel online reinforcement learning framework. We introduce a stable Identity-Preserving Preference Optimization that directly aligns the model across aesthetic quality, text-image alignment, and subject fidelity. To accurately measure subject fidelity, we introduce a Multi-ID Alignment Reward, which uses the Hungarian matching algorithm to maximize the overall match quality between multiple generated subjects and their references for precise scoring.
Experiments demonstrate that our approach, which decouples the learning of subject fidelity and human preference optimization into distinct training stages, achieves significant improvements over existing methods. Our main contributions are as follows:
@@
  • To our knowledge, we design the first online reinforcement learning framework tailored for multi-subject generation, introducing a Multi-ID Alignment Reward for stable and preference-aligned training.
  • We demonstrate improvements over existing methods, particularly in subject fidelity, achieving state-of-the-art performance in overall evaluations.
Figure 2: Visual comparison of attention maps. The ICL-based method UNO (left), fails to preserve subject fidelity. This is due to attention bleeding, where the double block’s attention regions for each subject are entangled, leading to attribute leakage. Our method overcomes this problem and maintains subject fidelity.
@@ 2 Related Work
Subject-driven Generation has attracted increasing attention [4, 19, 52, 49, 23, 20, 38, 39, 10, wang2024high, chen2024disenstudio]. These works can be broadly categorized into two types. (1) Fine-tuning-based methods [13, 48, 26]. These approaches achieve customization by fine-tuning part of the model’s parameters. (2) Tuning-free methods [65, 55, 31]. These approaches leverage large-scale training to eliminate the need for retraining when the subject changes. Recently, the powerful capabilities of foundation models based on the DiT [45] architecture have greatly enhanced the generation of multiple subjects, resulting in many excellent multi-subject generation works, e.g., HunyuanCustom [22], OmniControl [53], UniReal [5], UNO [59], OmniGen [62]. Some methods [liu2023customizable, 16, 56] address attribute leakage using user-provided spatial layouts to constrain attention during inference, which increases complexity for users. Our ICL-based method directly constrains attention between the reference and generated images during training, bypassing the text modality, enabling precise spatial alignment and requiring no extra user input at inference.
Reinforcement Learning for Text-to-Image Generation has become an active area of research. Initial strategies included policy gradient methods like Proximal Policy Optimization (PPO) [50, 2, 9, 18, 41, 70, xiao2025fastcomposer]. A subsequent major development is the adoption of Direct Preference Optimization (DPO) and its variants [54, 64, 66, 37, 68, 12, 29]. Some recent works introduce online RL technology, Group Relative Policy Optimization (GRPO)  [51], into Text-to-Image Generation, achieving significant performance gains. Flow-GRPO [35], DanceGRPO [63] introduce exploration by reformulating the deterministic Ordinary Differential Equation (ODE) of flow-matching models into a Stochastic Differential Equation (SDE). The improved MixGRPO [30] further boosts training efficiency with a mixed ODE-SDE framework. However, these methods have been limited to the basic text-to-image task, leaving the application of online reinforcement learning to multi-subject driven generation largely unexplored.
@@ 3 Preliminary
$\mathcal{L}_{diff}=\mathbb{E}_{t,z_{0},\epsilon\in\mathcal{N}(0,1)}\|v-v_{\theta}(z_{t},t,c_{text})\|^{2}.$ (1)
The DiT framework and Flow Matching are widely used in recent diffusion models, such as Stable Diffusion3 [8] and Flux [27]. In this paper, we use Flux as our base model.
Group Relative Policy Optimization (GRPO). GRPO struggles with ODE-based flow models because their deterministic systems lack the inherent stochasticity essential for reinforcement learning. So DanceGRPO [63] and Flow-GRPO [35] convert ODEs to SDEs, enabling stochastic reinforcement learning for image generation. MixGRPO [30] further improves efficiency by applying this only within a sliding time window during training. Given the prompt $c_{text}$, the training process in MixGRPO is similar to Flow-GRPO and DanceGRPO, but only optimizes the time steps sampled within the interval $S$. The behavior of this window is governed by key hyperparameters: the window size $w$ sets the number of consecutive timesteps to optimize at once, the shift interval $\tau$ determines how many training iterations pass before the window moves, and the window stride $s$ specifies how many timesteps the window advances during a shift. The final training objective is given by:
$J(\theta)=\mathbb{E}_{x\sim\pi_{\theta_{\text{old}}}(\cdot|c)}$ $\Bigg[\frac{1}{N}\sum_{i=1}^{N}\frac{1}{|S|}\sum_{t\in S}\min\Big(r_{i}^{t}(\theta)A_{i},$ (2)
$clip(r_{i}^{t}(\theta),1-\beta,1+\beta)A_{i}\Big)\Bigg],$
where $r_{i}^{t}(\theta)=\frac{\pi_{\theta}(x_{t+1}\mid x_{t},c)}{\pi_{\theta_{\text{old}}}(x_{t+1}\mid x_{t},c)}$ is the policy ratio and $A_{i}$ is the advantage score. $\beta$ is a hyperparameter that serves to clip the policy ratio, ensuring stable updates, and
$\quad A_{i}=\frac{R(x_{i}^{T},c)-mean\{R(x_{i}^{T},c)\}_{i=1}^{N}}{std\{R(x_{i}^{T},c)\}_{i=1}^{N}},$ (3)
where $R(x_{i}^{0},c)$ is provided by the reward model.
@@ 4 Methods
Figure 3: Overall pipeline of MultiCrafter. Our framework is built on two core innovations: (Top) Identity-Disentangled Attention Regularization uses positional supervision to prevent attribute leakage and the MoE-LORA architecture to boost model capacity for diverse scenarios; and (Bottom) the Identity-Preserving Preference Alignment framework employs a novel online reinforcement learning strategy with a Multi-ID Alignment Reward and the stable GSPO algorithm to align the model with human preferences.
In this section, we first explore the impact of highly coupled training objectives on the entire training process and present the overall architecture of our decoupled framework in Section˜4.1. Then, in Section˜4.2, we introduce Identity-Disentangled Attention Regularization, which we used during pre-training to achieve the goal of high subject fidelity. Finally, Section˜4.3 describes Identity-Preserving Preference Optimization (IPPO), which we used in the post-training stage to directly align the model with multi-dimensional human preferences.
4.1 Explore the In-Context-Learning
Given $N$ reference images of different subjects, we aim to generate images of these subjects according to a text prompt. Existing ICL-based methods encode these subjects into latent features, $\mathcal{Z}=\{\mathbf{z}_{i}^{ref}\}_{i=1}^{N}$, and leverage DiT-based attention mechanisms for customized generation. During training, these methods attempt to achieve high subject fidelity and conform to human preferences simultaneously, using a single reconstruction loss during the training phase. This forces the model to find a sub-optimal compromise between these two objectives, which especially reduces subject fidelity when subjects have similar attributes. Furthermore, this simple supervision method fails to adequately account for multi-dimensional human preferences. Therefore, we decouple the training process into pre-training and post-training stages, allowing the model to focus separately on subject fidelity and alignment with human preferences.
To explore how to improve subject fidelity during pre-training, we visualized the attention maps of a representative method [59], focusing on the double and single block structures. As shown on the left of Fig.˜2, we identify two key phenomena. First, the double blocks of Flux are far more pivotal in determining the spatial layout of the reference subject than the single block. Second, methods trained solely on a reconstruction loss lead to an undesired entanglement between subject-specific attention fields, causing attribute leakage and severely compromising subject fidelity. We can therefore infer a critical condition for high fidelity: the peak response within the double block’s attention scores, corresponding to a specific subject, must consistently align with that subject’s spatial region in the generated output. We leverage this in our method.
@@ 4.2 Identity-Disentangled Attention Regularization
Based on  Section˜4.1, the subject fidelity can be enhanced by strictly aligning the hot spot areas of the attention scores of double blocks with the spatial positions of the corresponding subjects in the generated image. We design a simple but effective regularization. Specifically, during the training process, for the latent space feature $\mathbf{z}_{i}^{ref}$ of the $i$-thth reference subject, we partition the reference feature $\mathbf{z}_{i}^{ref}$ into patches and apply positional encoding, resulting in a sequence of 1D tokens $\mathbf{z}_{i}^{r\prime}\in\mathbb{R}^{l\times c}$, $l$ is the number of tokens, $c$ is the number of channels. Then, we can obtain its attention map with the generated content $z_{t}$ at the $k$-th double block:
@@
where $\mathbf{Q}_{k,i}\in\mathbb{R}^{l\times c}$ is the query generated from the $i$-th subject image within the $k$-th double block, $\mathbf{K}\in\mathbb{R}^{l_{t}\times c}$ is the key produced by the noisy image latent tokens in the current layer, and $l_{t}$ denotes the number of tokens in the noisy image latent. For a model with $N$ double blocks, we can obtain the attention maps corresponding to the $i$-th subject from all blocks, which are aggregated into a set $\{m_{1}^{i},m_{2}^{i},...,m_{N}^{i}\}$. We then average and normalize this set to obtain the mean attention map $\hat{M}_{i}$. By pre-annotating the training data, we obtain the ground-truth mask ${M}_{i}$ corresponding to the $i$-th subject within the generated image. Notably, for human subjects, we exclusively use the facial region as the reference image and similarly focus only on the facial area for the generated image.
We employ the dice loss [42], a standard loss function for segmentation tasks, to minimize the discrepancy between each ground-truth mask ${M}_{i}$ and the corresponding mean attention map $\hat{M}_{i}$. The formulation is as follows:
@@
However, for efficient training, existing ICL-based methods usually use LoRA to fine-tune the model. But the significant variance in the spatial layouts generated by different prompts and subjects poses a challenge for a standard LoRA, whose limited capacity is insufficient to master the complex task. To address this, and inspired by the success of MoE-LoRA in multitask tuning [11, 36, 15], we adopt an MoE-LoRA architecture to expand model capacity. This allows different experts to focus on the spatial layout of various scenarios, effectively alleviating this problem. We strategically integrate the MoE-LoRA into the output feed-forward network (FFN) layers of the Flux, while other layers are adapted using standard LoRA for parameter efficiency. Since the use of experts in this situation is difficult to define manually, we choose to let the model learn this implicitly. Specifically, given an input vector $h$ to the FFN layer, we define the number of experts as $N_{e}$, the rank of each LoRA as $r$, and the scaling factor as $\alpha$. A lightweight gating network, $g_{\theta}$, which dynamically routes the input vector $h$ to the most suitable experts, computes a vector of logits, $p\in\mathbb{R}^{N_{e}}$, for the experts:
$p=\operatorname{Softmax}(\operatorname{TopK}(W_{g}\cdot h,k)),$ (7)
@@
$h_{out}=\text{FFN}(h)+\sum_{i=1}^{N_{e}}p_{i}\cdot\left(\frac{\alpha}{r}\cdot W_{B}^{i}\cdot W_{A}^{i}\cdot h\right).$ (8)